OpenStudy (waheguru):

what is teh equation if m=4/3 and the coordinates are (-2,0) HELPPPPPPPPPPPPPP what is teh equation if m=4/3 and the coordinates are (-2,0) HELPPPPPPPPPPPPPP @Mathematics

hero (hero):

y - y_1 = m(x-x_1) y - 0 = (4/3)(x+2) y = (4x+8)/3 y = 4/3x + 8/3